### v1

@@ I Introduction
However, the aforementioned methods for SPSs rely entirely or partially on implicit numerical methods. This is because the fast dynamics introduced by small time constants require indispensable numerical stability when using step sizes much larger than the time constants [12]. Common explicit numerical methods, such as Forward Euler and the Runge-Kutta methods, often require prohibitively small time steps to stabilize the fast dynamics, leading to increased computational cost and inefficiency [13]. Meanwhile, implicit methods, although more stable for SPSs, involve solving nonlinear equations at each iteration, which can be computationally intensive and may become inefficient for large-scale or highly nonlinear systems [14]. IMEX methods can take advantage of the efficiency of explicit methods and the stability of implicit methods; however, they require an explicit separation of the fast and slow dynamics which is not always simple or trivial.
@@
An important observation is that SPSs can be usually be related to differential-algebraic equations (DAEs) when the perturbation parameter approaches zero [13]. In this limit, the fast dynamics effectively become algebraic constraints, and the system exhibits characteristics similar to DAEs. Numerical methods for solving DAEs often employ variable-step or multirate strategies to efficiently handle the coupled differential and algebraic components without resorting to excessively small time steps [15]. Inspired by these techniques, SMES incorporates a multirate approach with variable time steps which stabilizes classical explicit numerical methods to handle the fast dynamics. This approach maintains the simplicity and ease of implementation of explicit methods while enhancing computational efficiency and providing numerical stability guarantees. Moreover, since SMES is based on explicit schemes, it is better suited for systems with strong nonlinearity compared to implicit methods.
The remainder of this paper is organized as follows: In Section II, we provide the relevant background on SPSs. In Section III, we present the limitation of classical explicit methods on SPSs using the Forward Euler method. In Section IV, we introduce the Stabilized Multirate Explicit Scheme (SMES) based on the Forward Euler method, outlining its theoretical foundation and implementation details. In Sections V and VI, we perform stability and error analyses of the proposed method. In Section VII, we present numerical illustrations to demonstrate the efficiency of our method on SPSs. Finally, Section VIII concludes the paper.
I-A Notation
If $f$ and $g$ are real-valued functions defined on a subset of $\mathbb{R}_{0}^{+}$ including a neighborhood of 00, we write $f=O(g)$ to mean that there are positive constants $K$ and $\delta$ such that
@@ II Preliminary
Singularly perturbed systems (SPSs) are dynamical systems characterized by multiple time scales. These systems are generally described by the state-space form [16]
@@
$0=g(\bar{x},\bar{z},\bar{u},0,t).$ (3)
Eq. (2) is said to be in standard form if and only if the following assumption concerning (3) is satisfied:
Assumption 1. In a domain of interest, (3) has $k\geqslant 1$ distinct real roots [5]
@@
This assumption ensures a well-defined $n$-dimensional reduced model of the full system in (2), described as the differential-algebraic equation (DAE)
@@
$\bar{z}$ $=\bar{\varphi}(\bar{x},\bar{u},t).$
Eq. (5) is also called a quasi-steady-state form of (2) because the fast dynamics in $z$, whose rate $\dot{z}=g/\varepsilon$ can be large when $\varepsilon$ is small, may rapidly converge to the equilibrium manifold $\mathcal{M}_{0}\colon\bar{z}=\bar{\varphi}(\bar{x},\bar{u},t)$.
@@
Remark 1. In this paper, we consider the more general case where the system may not be in standard form, or may not have an explicit separation of fast and slow dynamics as in (2), which is not always straightforward to determine. Nevertheless, to simplify the subsequent analysis, we consider the standard form of singularly perturbed systems that satisfies Assumption 1.
@@ III Forward Euler Method
$\dot{x}$ $=A_{11}x+A_{12}z,\quad x(t_{0})=x^{0},$ (7)
$\varepsilon\dot{z}$ $=A_{21}x+A_{22}z,\quad z(t_{0})=z^{0},$
where $x\in\mathbb{R}^{n}$ and $z\in\mathbb{R}^{m}$. If $A_{22}$ is nonsingular, there exists a matrix $L(\varepsilon)\in\mathbb{R}^{m\times n}$ satisfying
@@
Accordingly, the system in (7) can be rewritten in block-triangular form, to separate the slow and fast dynamics, as
@@ IV Stabilized Multirate Forward Euler Method
Given the limitations of classical explicit numerical methods, we introduce a Stabilized Multirate Explicit Scheme (SMES) that handles singular perturbations without requiring prohibitively small step sizes, thus allowing for faster and stable simulations of SPSs using explicit methods. In this section, we illustrate how SMES works by augmenting the FEM with the proposed scheme, resulting in a stabilized multirate Forward Euler (SMFE) method.
@@
Note that the total time interval covered by these $N+1$ steps is $\Delta$. This stabilized multirate Forward Euler (SMFE) method is described in Algorithm 1.
@@ V Stability analysis
For small $\varepsilon$ but fixed $\Delta$ and $N$, condition (14) yields
$|1+\Delta\hat{\lambda}_{\text{slow}}|<1,$ (16)
where we used the approximation $(1+\delta)^{N}\approx 1+N\delta$ for small $\delta$ and neglected the $O(\varepsilon)$ terms. Therefore, for the slow subsystem, the stability condition for the Forward Euler discretization is recovered.
Next, we consider the fastest eigenvalue $\hat{\lambda}_{\text{fast}}$ as in (10) and define the scaled eigenvalue $\tilde{\lambda}=\varepsilon\hat{\lambda}_{\text{fast}}$. For $\varepsilon\ll 1$, condition (14) is approximately equivalent to
$N\geqslant-\frac{\ln|1+\Delta\hat{\lambda}_{\text{fast}}-N\Delta\tilde{\lambda}|}{\ln|1+\Delta\tilde{\lambda}|}.$
@@
$N\geqslant-\frac{\ln(\Delta|\tilde{\lambda}|/\varepsilon)}{\ln|1+\Delta\tilde{\lambda}|}=O\left(\ln\left(\frac{1}{\varepsilon}\right)\right).$ (17)
@@ VI Error analysis
$\eta(t+N\Delta\varepsilon)$ $\leqslant|1+\Delta\varepsilon\hat{\lambda}_{\text{fast}}|^{-\frac{\ln(\Delta|\tilde{\lambda}|/\varepsilon)}{\ln|1+\Delta\tilde{\lambda}|}}\eta_{\Delta}.$ (18)
Since $\tilde{\lambda}=\varepsilon\hat{\lambda}_{\text{fast}}$, (18) simplifies to
$\eta(t+N\Delta\varepsilon)$ $\leqslant\frac{\varepsilon\eta_{\Delta}}{\Delta|\tilde{\lambda}|}=O(\varepsilon).$ (19)
@@
Since the $N$ small-step iterations introduce an error of $O(\varepsilon)$, the error from the Forward Euler discretization in (20), which is $O(\Delta^{2})$ [13], becomes the dominant source when $O(\Delta^{2})\gg O(\varepsilon)$. Therefore, the error introduced by the SMFE method on the full model over the time interval $\Delta$ is very close to the discretization error produced during the simulation of one step length $\Delta(1-N\varepsilon)$ on the reduced system.
@@ VII-A Efficiency of the SMFE method
Iterations $5.00\times 10^{6}$ $1.78\times 10^{3}$ $3.53\times 10^{3}$ $2.80\times 10^{4}$ $7.05\times 10^{3}$ $5.60\times 10^{4}$ $5.61\times 10^{5}$

### v2

@@ I Introduction
However, the aforementioned methods for SPSs rely entirely or partially on implicit numerical methods. This is because the fast dynamics introduced by small time constants require indispensable numerical stability when using step sizes much larger than the time constants [12]. Common explicit numerical methods, such as Forward Euler and the explicit Runge-Kutta methods, often require prohibitively small time steps to stabilize the fast dynamics, leading to increased computational cost and inefficiency [13]. Meanwhile, implicit methods, although more stable for SPSs, involve solving nonlinear equations at each iteration, which can be computationally intensive and may become inefficient for large-scale or highly nonlinear systems [14]. IMEX methods can take advantage of the efficiency of explicit methods and the stability of implicit methods; however, they require an explicit separation of the fast and slow dynamics, which is not always simple or trivial.
@@
An important observation is that SPSs can usually be related to differential-algebraic equations (DAEs) when the perturbation parameter approaches zero [13]. In this limit, the fast dynamics effectively become algebraic constraints, and the system exhibits characteristics similar to DAEs. Numerical methods for solving DAEs often employ variable-step or multirate strategies to efficiently handle the coupled differential and algebraic components without resorting to excessively small time steps [15]. Inspired by these techniques, SMES incorporates a multirate approach with variable time steps, which stabilizes classical explicit numerical methods to handle the fast dynamics. This approach maintains the simplicity and ease of implementation of explicit methods while enhancing computational efficiency and providing numerical stability guarantees. Moreover, since SMES is based on explicit schemes, it is better suited for systems with strong nonlinearity compared to implicit methods.
The remainder of this paper is organized as follows: In Section II, we provide the relevant background on SPSs. In Section III, we present the limitation of classical explicit methods on SPSs using the Forward Euler method. In Section IV, we introduce the SMES based on the Forward Euler method, outlining its theoretical foundation and implementation details. In Sections V and VI, we perform stability and error analyses of the proposed method. In Section VII, we present numerical illustrations to demonstrate the efficiency of our method on SPSs. Finally, Section VIII concludes the paper.
Notation. If $f$ and $g$ are real-valued functions defined on a subset of $\mathbb{R}_{0}^{+}$ including a neighborhood of 00, we write $f=O(g)$ to mean that there are positive constants $K$ and $\delta$ such that
@@ II Preliminary
SPSs are dynamical systems characterized by multiple time scales. These systems are generally described by the state-space form [16]
@@
$0=g(\bar{x},\bar{z},\bar{u},0,t),$ (3)
where the bar is used to indicate that the variables belong to a system with $\varepsilon=0$. Eq. (2) is said to be in standard form if and only if the following assumption concerning (3) is satisfied:
Assumption 1. [5] In a domain of interest, (3) has $k\geqslant 1$ distinct real roots
@@
This assumption ensures a well-defined $n$-dimensional reduced model of the full system in (2) corresponding to each root, described as the differential-algebraic equation (DAE)
@@
$\bar{z}$ $=\bar{\varphi}_{i}(\bar{x},\bar{u},t).$
Eq. (5) is also called a quasi-steady-state form of (2) because the fast dynamics in $z$, whose rate $\dot{z}=g/\varepsilon$ can be large when $\varepsilon$ is small, may rapidly converge to the equilibrium manifold $\mathcal{M}_{0}\colon\bar{z}=\bar{\varphi}_{i}(\bar{x},\bar{u},t)$.
@@
Remark 1. In this paper, we consider the more general case where the system may not be in standard form, or may not have an explicit separation of fast and slow dynamics as in (2), which is not always straightforward to determine. Nevertheless, to simplify the subsequent analysis, we consider the standard form of singularly perturbed systems that satisfy Assumption 1.
@@ III Forward Euler Method
$\dot{x}(t)$ $=A_{11}x(t)+A_{12}z(t),\quad x(t_{0})=x^{0},$ (7)
$\varepsilon\dot{z}(t)$ $=A_{21}x(t)+A_{22}z(t),\quad z(t_{0})=z^{0},$
where $x(t)\in\mathbb{R}^{n}$, $z(t)\in\mathbb{R}^{m}$, and the time variable $t$ is explicitly written out to facilitate subsequent analysis. If $A_{22}$ is nonsingular, there exists a matrix $L(\varepsilon)\in\mathbb{R}^{m\times n}$ satisfying
@@
The system in (7) can then be rewritten in block-triangular form [5], to separate the slow and fast dynamics, as
@@ IV Stabilized Multirate Forward Euler Method
Given the limitations of classical explicit numerical methods, we introduce a SMES that handles singular perturbations without requiring prohibitively small step sizes, thus allowing for faster and stable simulations of SPSs using explicit methods. In this section, we illustrate how SMES works by augmenting the FEM with the proposed scheme, resulting in a stabilized multirate Forward Euler (SMFE) method.
@@
Note that the total time interval covered by these $N+1$ steps is $\Delta$. This SMFE method is described in Algorithm 1.
@@ V Stability analysis
For small $\varepsilon$ but fixed $\Delta$ and $N$, we use the approximation $(1+\Delta\varepsilon\lambda)^{N}\approx 1+N\Delta\varepsilon\lambda$ and neglected the $O(\varepsilon)$ terms, thus condition (14) yields
$|1+\Delta\hat{\lambda}_{\text{slow}}|<1.$ (16)
Therefore, for the slow subsystem, the stability condition for the Forward Euler discretization is recovered.
Next, we consider the fastest eigenvalue $\hat{\lambda}_{\text{fast}}$ as in (10) and define the scaled eigenvalue $\tilde{\lambda}=\varepsilon\hat{\lambda}_{\text{fast}}=O(1)$. For $\varepsilon\ll 1$, condition (14) is equivalent to
$N>-\frac{\ln|1+\Delta\hat{\lambda}_{\text{fast}}-N\Delta\tilde{\lambda}|}{\ln|1+\Delta\tilde{\lambda}|}.$
@@
$N>-\frac{\ln(\Delta|\tilde{\lambda}|/\varepsilon)}{\ln|1+\Delta\tilde{\lambda}|}=O\left(\ln\left(\frac{1}{\varepsilon}\right)\right).$ (17)
@@ VI Error analysis
$\eta(t+N\Delta\varepsilon)$ $<|1+\Delta\varepsilon\hat{\lambda}_{\text{fast}}|^{-\frac{\ln(\Delta|\tilde{\lambda}|/\varepsilon)}{\ln|1+\Delta\tilde{\lambda}|}}\eta_{\Delta}.$ (18)
Given $\tilde{\lambda}=\varepsilon\hat{\lambda}_{\text{fast}}$ and the principle $a^{b}=e^{b\ln a}$ for positive $a$, (18) can be simplified to
$\eta(t+N\Delta\varepsilon)$ $<\frac{\varepsilon\eta_{\Delta}}{\Delta|\tilde{\lambda}|}=O(\varepsilon).$ (19)
@@
Since the $N$ small-step iterations introduce an error of $O(\varepsilon)$, the error from the Forward Euler discretization in (20), which is $O(\Delta^{2})$ [13], becomes the dominant source when $O(\Delta^{2})\gg O(\varepsilon)$. Therefore, the error introduced by the SMFE method on the full model over time interval $\Delta$ is very close to the discretization error produced during the simulation of step length $\Delta(1-N\varepsilon)$ on the reduced system.
@@ VII-A Efficiency of the SMFE method
Iterations $5,000,000$ $1,780$ $3,530$ $28,000$ $7,050$ $56,000$ $561,000$